# thanks: ISTI Fellow

### v1

@@ Appendix A Interpreting the DiCo20 $\tilde{\Lambda}$ Results
The DiCo20 $\tilde{\Lambda}$ and $R_{\rm{1.4}}$ posteriors in Figures 2 and 5 are appreciably different from those corresponding to the KruFo20 and Nedora21 ejecta fits, favoring much larger radii and $\tilde{\Lambda}$ than informed by the GW likelihood. We examine the DiCo20 ejecta contours in Figure 8. The velocity prescription remains the same, while the disk mass and dynamical ejecta fits change. The dynamical ejecta contours behave similarly to the KruFo20 contours, with larger values of $\delta$ required to produce a higher quantity of dynamical ejecta. The disk mass contours (dashed lines) are the most different between the two fits. Whereas in the KruFo20 fits, disk mass was weakly dependent on $\delta$ and primarily set by $R_{\rm{1.4}}$, the DiCo20 fits behave oppositely, with disk mass defined by $\delta$ and completely agnostic to $R_{\rm{1.4}}$. This reversed dependency enables the DiCo20 posteriors to find solutions near $\delta=0$ with low $\chi_{\rm{eff}}$, in line with the GW-informed likelihood; however, the lack of a $R_{\rm{1.4}}$ constraint results in the uncharacteristically large $\tilde{\Lambda}$ values observed in Figure 2.

### v2

@@ Appendix A Interpreting the DiCo20 $\tilde{\Lambda}$ Results
The DiCo20 $\tilde{\Lambda}$ and $R_{\rm{1.4}}$ posteriors in Figures 2 and 3 are appreciably different from those corresponding to the KruFo20 and Nedora21 ejecta fits, favoring much larger radii and $\tilde{\Lambda}$ than informed by the GW likelihood. We examine the DiCo20 ejecta contours in Figure 8. The velocity prescription remains the same, while the disk mass and dynamical ejecta fits change. The dynamical ejecta contours behave similarly to the KruFo20 contours, with larger values of $\delta$ required to produce a higher quantity of dynamical ejecta. The disk mass contours (dashed lines) are the most different between the two fits. Whereas in the KruFo20 fits, disk mass was weakly dependent on $\delta$ and primarily set by $R_{\rm{1.4}}$, the DiCo20 fits behave oppositely, with disk mass defined by $\delta$ and completely agnostic to $R_{\rm{1.4}}$. This reversed dependency enables the DiCo20 posteriors to find solutions near $\delta=0$ with low $\chi_{\rm{eff}}$, in line with the GW-informed likelihood; however, the lack of a $R_{\rm{1.4}}$ constraint results in the uncharacteristically large $\tilde{\Lambda}$ values observed in Figure 2.